OpenStudy (anonymous):

Solve for x in the following proportion: x over 5 equals 3 over 2 A. x = 10 B. x = 15 C. x = 15/2 D. x = 15/3

OpenStudy (anonymous):

x/5 = 3/2

OpenStudy (anonymous):

c

OpenStudy (anonymous):

it is also simple just cross multiply x/5 = 3/2 2*x = 5*3 2x = 15 x = 15/2 x = 7.5

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

anymore questions

OpenStudy (anonymous):

two more i'm about to post .